Hank, a calendar-year taxpayer, uses the cash method of accounting for his sole proprietorship. In late December, he performed $44,000 of legal services for a client. Hank typically requires his clients to pay his bills immediately upon receipt. Assume his marginal tax rate is 32 percent this year and will be 37 percent next year, and that he can earn an after-tax rate of return of 9 percent on his investments.

a. What is the after-tax income if Hank sends his client the bill in December?

b. What is the after-tax income if Hank sends his client the bill in January?
Answer #1

a.After tax income = Income before tax(1-Tax rate)

= 44,000*(1-32%)

= $29,920

b.After tax income in January = 44,000(1-37%) = $27,720